OpenStudy (anonymous):

Find the volume of a cone with radius 6 m and height 20 m. Choose one answer. a. 720π m^3 b. 240π m^3 c. 120π m^3 d. 40π m^3

OpenStudy (anonymous):

@jim_thompson5910 i got 753 and i did all the steps but no right answer... help me?

jimthompson5910 (jim_thompson5910):

Area of the base = ???

OpenStudy (phi):

Did you use \[ V= \frac{1}{3} \pi r^2 h \]

OpenStudy (anonymous):

@jim_thompson5910 sorry i had to run some errands for my mom but the area of the base is 113.04

OpenStudy (anonymous):

@jim_thompson5910 am i right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now multiply that with the height

OpenStudy (anonymous):

2260.8

jimthompson5910 (jim_thompson5910):

then divide by 3

OpenStudy (anonymous):

i still got 753.6!

OpenStudy (anonymous):

thats wrong!

jimthompson5910 (jim_thompson5910):

ok which matches to which answer

jimthompson5910 (jim_thompson5910):

i know, each answer has a pi in it

jimthompson5910 (jim_thompson5910):

one of those answers is roughly equal to 753.6

OpenStudy (anonymous):

720

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

no, try doing the calculations without 3.14 and use pi instead

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i really dont know i did and i got 753.98223687 this time

OpenStudy (phi):

when you find the area using \[ \pi r^2\] leave pi in. do not use 3.14 you get \[ 36 \pi\]

OpenStudy (anonymous):

i dont have my graphing calc right now so im using a scientific one right now

OpenStudy (anonymous):

but ohk yes i got 36

OpenStudy (phi):

V= 1/3 pi r^2 h if you leave the pi to one side: \[ \pi (\frac{1}{3} 36 \cdot h )\] just figure out the stuff in the parens

OpenStudy (anonymous):

i got 240 woowww how'd yo do that... how was i getting 753?

OpenStudy (anonymous):

I got B for this is that right? Find the volume of a square pyramid with base edges 4 ft. and height 15 ft. Choose one answer. a. 20 cu. ft b. 80 cu. ft c. 240 cu. ft d. 60 cu. ft

OpenStudy (phi):

the answer is \[ 240 \cdot \pi \] 240 times pi if you replace pi with 3.14159... you get 753. with lots of decimals people just write it 240 pi and if someone wants the exact number, they can multiply it out themselves.

OpenStudy (phi):

Yes to b. 80 cu. ft

OpenStudy (anonymous):

oh ohk thanks a lot can you check my answer above?

OpenStudy (anonymous):

thanks :)